The subatomic particle called the Delta (1232) is
Answer #1

a) The uncertainity relation between energy and time is ∆t*∆E ∼ h/4pi = 5.27*10^-35

Then   ∆E = 5.27*10^-35/10^-23 = 5.27*10^-12 J

b) The uncertainity relation between momentum and position is ∆x*∆p ∼ h/4pi = 5.27*10^-35

  Then   ∆p = 5.27*10^-35/5*10^-15 = 1.054*10^-20

Then minimum K.E . = (∆p)^2/2m = (1.054*10^-20 )^2/(2*1.67*10^-27) = 3.33*10^-14 J